# Teaching Large Language Models to Regress Accurate Image Quality Scores Using Score Distribution

### v2

@@ Appendix D Extensions
Score regression helps quality description. We investigate whether score regression tasks can enhance general low-level perception tasks, i.e., language-based quality description tasks. The low-level perception tasks are evaluated on Q-Bench [68]. For score regression, we adopt the same training techniques as in the main paper, and train the model on the KonIQ, SPAQ, and KADID datasets. The low-level perception tasks are trained using the next token prediction paradigm, with Q-Instruct [69] as the training dataset. Q-Instruct [69] has shown that co-training with or pre-training on high-level tasks can improve performance on low-level perception tasks. Similarly, we evaluate two strategies: (a) co-training with score regression tasks, and (b) pre-training on score regression tasks. Some questions in the Q-Instruct dataset are quite similar to the questions of score regression, which may confuse the model. Therefore, we append the questions of score regression with a specific suffix, “Answer the question with levels.”, to specify the task.

### v3

@@ Appendix D Extensions
Score regression helps quality description. We investigate whether score regression tasks can enhance general low-level perception tasks, i.e., language-based quality description tasks. The low-level perception tasks are evaluated on Q-Bench [qbench]. For score regression, we adopt the same training techniques as in the main paper, and train the model on the KonIQ, SPAQ, and KADID datasets. The low-level perception tasks are trained using the next token prediction paradigm, with Q-Instruct [qinstruct] as the training dataset. Q-Instruct [qinstruct] has shown that co-training with or pre-training on high-level tasks can improve performance on low-level perception tasks. Similarly, we evaluate two strategies: (a) co-training with score regression tasks, and (b) pre-training on score regression tasks. Some questions in the Q-Instruct dataset are quite similar to the questions of score regression, which may confuse the model. Therefore, we append the questions of score regression with a specific suffix, “Answer the question with levels.”, to specify the task.